OpenStudy (anonymous):

find the point of intersection of the two graphs plot the graphs of both equations on the same coordinate plane.find and label the points of intersection of the two graphs Y=-2X+3 Y=-2(X-4)^2

Parth (parthkohli):

This is merely a system of equations.\[-2x + 3= -2(x - 4)^2\]by substitution

OpenStudy (anonymous):

equate two functions

OpenStudy (anonymous):

parthkohli and how to change to general form?

OpenStudy (anonymous):

is this correct -2x+3=-2(x-4)^2 -2x^2+18x-13=0

OpenStudy (anonymous):

probably not, try again

OpenStudy (anonymous):

-2x^2+18x-35=0

OpenStudy (anonymous):

so it is a quadratic equation,and we solve by the root formula right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@lissafi , note that when you equate two functions, you can find x intersections as well as y intersections

OpenStudy (anonymous):

here above, you will find x intersections, to find y, reformulate each equation so that it starts with x=...., but, beware there you have a square, so you can lose information by reformulation

OpenStudy (anonymous):

what is the square root of 386?

OpenStudy (anonymous):

@lissafi, do you have a calculator?

OpenStudy (anonymous):

is too hard for me

OpenStudy (anonymous):

x=(-18+/-root(604))/-4

OpenStudy (anonymous):

+/- means that: + or -

OpenStudy (anonymous):

not "over"

OpenStudy (anonymous):

Its 19.64